A man in a balloon rising with constant velocity of 6 m/s propels a ball upwards with a velocity of 9 m/s, relative to the balloon. After what time interval will be the bal return to the balloon g-=9.8 m/s?

Grade:12th pass

2 Answers

Arun
5 years ago
Speed of balloon w.r.t ground=6 m/s (Upwards)
Speed of stone w.r.t balloon=9 m/s
So, speed of stone w.r.t ground=6 +9=15 m/s
​u=15 m/s (upwards)
Using relation
v=u+gt
=15-10 x t = 0
 
t = 15/10 = 1.5 sec
BABU S
3 years ago
Man with balloon raising velocity = 6 m/s (constant velocity)
ball upward velocity = 9 m/s 
man with balloon covering vertical distance= 6t
distance of ball reached the balloon= 9t-(1/2*9.81 t2)
both distance are equal
6t=9t-(1/2*9.81 t2)
t=0.61 s